# Edge-Spreading Raptor-Like LDPC Codes for 6G Wireless Systems

### v1

@@ IV-A2 Spreading Optimization over the Unified Graph
Unlike [25, 24], there is no need to repeatedly run the above message-passing algorithm on $\mathcal{G}_{\text{U}}$ to update the cycle distribution when changes occur in $\mathcal{G}_{\text{U}}$. Instead, we can simply adjust the polynomial $p_{v}$ in (13) to measure the impact of reallocating a specific edge on the cycle distribution. It is noteworthy that the extra complexity of such an optimization on $\mathcal{G}_{\text{U}}$ is negligible. Specifically, if the edge $e_{v}^{c}$ is re-assigned from $\mathbf{T}_{c,v}$ to ${\widetilde{\mathbf{T}}_{c,v}}$ (where ${\widetilde{\mathbf{T}}_{c,v}}\in[0,\omega]$ and ${\widetilde{\mathbf{T}}_{c,v}}\neq\mathbf{T}_{c,v}$), let $\theta={\widetilde{\mathbf{T}}_{c,v}}-\mathbf{T}_{c,v}$, we modify the original $p_{v}$ by two steps: 1 Add $\theta$ to all accumulated path values $\phi$ of the initial $\gamma_{c}$ variable in the polynomial $p_{v}$ 2 Subtract $\theta$ from the $\phi$ of all $\gamma$ variables in the monomial passed to the VN $v$ along the edge $e_{c}^{v}$ (as both $e_{v}^{c}$ and $e_{c}^{v}$ belong to the same undirected edge, they are equivalent in the spreading process).
@@ IV-C Design Flow of ESRL Codes
The HRC design process is outlined in Algorithm 2. In Step 1, under the ESRL code constraints from Section III-B, DE is used to determine the optimal distribution of $d_{v}$, denoted as the polynomial $p_{d_{v}}^{\ast}$. Step 2 employs the progressive edge growth (PEG) algorithm [46] and the approximate cycle extrinsic message degree (ACE) constraint [47] to maximize local girth and to ensure the graph connectivity of short cycles on $\mathbf{B}_{\text{HRC}}$, respectively, which are two classical methods for designing finite-length LDPC codes. Notably, threshold analysis on $\mathbf{B}_{\text{HRC}}$ is not part of our flow, aligning with the objective of ESRL codes in Section III-A to focus on constructing a better coupled protomatrix $\mathbf{B}_{\text{SC}}$. The next step is to use the unified graph $\mathcal{G}_{\text{U}}$ in Section IV-A and the modified coupled RCA in Section IV-B to optimize $\mathbf{B}_{\text{SC}}$. In Step 3, the matrix $\mathbf{T}$ is initially set to $\mathbf{0}$, indicating that all edges start in $\mathbf{B}_{0}$. The unified graph $\mathcal{G}_{\text{U}}$ is then employed to optimize the spreading strategy, reallocating each edge to minimize cycle count. In Step 4, $\mathbf{B}_{\text{SC}}$ is constructed using the code profile $\mathbf{B}$$\mathbf{T}$, and $\mathbf{Q}$, followed by modified coupled RCA to compute the threshold. Since $\mathbf{Q}$ has a fixed structure without additional design, it directly participates in the threshold analysis of $\mathbf{B}_{\text{SC}}$. Finally, after $I_{\text{HRC}}$ iterations, the optimal $\mathbf{B}_{\text{HRC}}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold are selected.
@@ 6 Step 3: Optimize over unified graph $\mathcal{G}_{\text{U}}$
7    $\mathbf{T}\leftarrow\mathbf{0}$; //initialize as $\mathbf{B}_{0}$
@@ 14 Repeat Step 2 to Step 4 $I_{\text{HRC}}$ times;
The IME design process is summarized in Algorithm 3. Compared to the overall design of the HRC in Algorithm 2, the IME is constructed row-wise until the full matrix of size $m\times n$ is completed. In the ESRL code, we find that the performance of the growing IME matrix is highly sensitive to the previously designed sections, especially after edge spreading, i.e., overemphasizing the performance of a specific intermediate rate will lead to significant performance degradation in subsequent matrix extensions. To address this issue, instead of using a greedy approach to explore a local optimum for each newly added row as in [17], we start with a function called ${\small\textsf{GlobalWeightRef}}(\cdot)$ in Step 1, which essentially runs Algorithm 2 to design a reference $\mathbf{B}$ of size $m\times n$ directly and thus offers a preliminary IME example. Note that although this reference is not optimal, it can provide a rough weight guideline $\mu_{\text{IME}}$ for the IME design to avoid being trapped in local optima for intermediate rows. Similarly, Step 2 involves using the PEG and ACE algorithms for the newly added row. In Step 3, the unified graph $\mathcal{G}_{\text{U}}$ is generated to optimize the spreading strategy. It is noteworthy that when designing the $c$-th row of the IME, the edges and corresponding allocations of previous $c-1$ rows must remain unchanged to preserve the optimization results obtained before. Besides, certain special structures in ESRL codes, such as assigning the diagonal of the SSC to $\mathbf{B}_{0}$, need to be retained. Finally, we run the process $I_{\text{IME}}$-times for each row to search for optimal $\mathbf{B}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold, until completing the entire ESRL protomatrix design $\mathbf{B}_{\text{SC}}$.
@@ 2    $\mathbf{B}^{\ast}\leftarrow\mathbf{B}_{\text{HRC}}^{\ast}$;
4    $\mu_{\text{IME}}\leftarrow{\small\textsf{GlobalWeightRef}}(\mathbf{B}_{\text{HRC}},m,n)$;
@@ 5 Step 2: Generate IME row by row
8          $\mathbf{B}\leftarrow\mathsf{PEG}(\mathbf{B}^{\ast},\mu_{\text{IME}})$;
@@ V Puncturing Scheme and Decoding Algorithm
In this section, we introduce the puncturing scheme and high-throughput SLME decoding tailored to ESRL codes.
@@ V-A Puncturing Scheme
Figure 8: Puncturing scheme of the ESRL code, where the proposed semi-layered multi-engine decoding is running.
@@
V-B SLME Decoding of ESRL codes
@@
In this section, we propose a high-throughput semi-layered decoding algorithm using a multi-engine architecture in [12], tailored to ESRL codes. This algorithm can be flexibly configured as either high-performance BP decoding of the entire $\mathbf{H}_{\text{SC}}$ or as low-complexity windowed BP sliding along the diagonal of $\mathbf{H}_{\text{SC}}$. Let $W$ denote the window size, $S$ the sliding step, and $I_{\text{max}}$ the maximum number of decoding iterations within the window at each step. For instance, by setting $W\!=\!S\!=\!L+\omega$, SLME decoding corresponds to traditional BP decoding with $I_{\text{max}}$ for ESRL codes.
Figure 10: ACE spectrum of the protograph example of ESRL codes.
@@
While coupling generally limits parallel decoding, ESRL codes without parallel edges can perform semi-layered decoding through a more compact multi-engine architecture compared to [12], achieving higher parallelism. Specifically, we apply individual layered decoding to each $\mathbf{U}$ (defined in (2) of Section III-A and also shown in Fig. 3), without causing memory read-write conflicts. In Fig. 8, with a window size of $W=4$, conventional layered decoding proceeds row-wise from top to bottom (indicated by a grey arrow). In contrast, SLME decoding can deploy four identical sub-decoders working in parallel (each running layered decoding shown by a light red arrow in Fig. 8), thus resulting in a 4×4$\times$4 × throughput gain in the example.
@@ VI-A Design Example of ESRL Codes
Fig. 9 illustrates an example ESRL protomatrix family designed for $k=22$ (in each batch). We intentionally set $k$ equal to that of 5G-NR to facilitate fair comparisons in Section VI-B. Following SC-LDPC decoder implementations [41], we set $\omega=1$ to simplify the code design. Moreover, the coupling length $L$ is set as $10$ to balance the rate loss and the underlying blocklength for a given frame length. Let $m^{\prime}=4$$n^{\prime}=26$$m=40$, and $n=62$. Hence, the HRC protomatrix has dimensions of $4\times 26$. The IME consists of the IRC protomatrix (sized $36\times 26$) and the SSC protomatrix (sized $36\times 36$). Considering that the first VN in the HRC is always punctured ($\rho=1$), the example ESRL code can support $37$ distinct rates, ranging from the highest rate of $0.87$ to the lowest rate of $0.34$. The uncoupled protomatrix $\mathbf{B}$ has a size of $40\times 62$ and a weight of $314$. Note that with this configuration we mark six classical rates in Fig. 9: $0.87~{}(4\times 26)$$0.8~{}(6\times 28)$$0.72~{}(9\times 31)$$0.63~{}(13\times 35)$$0.52~{}(20\times 42)$, and $0.34~{}(40\times 62)$, which are also supported by 5G-NR and are used for the subsequent comparisons. For instance, the highest rate is computed as $R=(22\times 10)/(26\times 10+4-1\times(10+1))\approx 0.87$ using (5). The edge-spreading matrix $\mathbf{T}$ is also reported in Fig. 9, with $\mathbf{B}_{0}$ and $\mathbf{B}_{1}$ distinguished by colors. As mentioned in Section III-C, the tail matrix $\mathbf{Q}$ has a fixed tri-diagonal structure, marked as black in Fig. 9, since it does not participate in the spreading process. In terms of connectivity of short cycles, in the lowest rate protograph, every cycle of length up to $8$ is connected to at least $12$ extrinsic CNs. For $4$-cycles, the value of ACE $\eta$ ranges from $6$ to $12$. Fig. 10 also shows the ACE spectrum of $6$-cycles and $8$-cycles. It is noteworthy that given these ACE values are obtained from the protograph before lifting, the actual graph connectivity will be even better in $\mathbf{H}_{\text{SC}}$ of the ESRL code after lifting.
Fig. 11 shows the frame error rate (FER) performance of the example ESRL codes under different lifting sizes. Let $\mathcal{C}$ define a series of rate-compatible LDPC codes with $K$ information bits, and $\mathcal{C}(N,K)$ be one of its code patterns with a rate $R=K/N$. Based on $Z\in\{39,78,390\}$, three series of ESRL codes are defined: $\mathcal{C}_{\text{SC}}^{\text{I}}$ $(K=8580)$$\mathcal{C}_{\text{SC}}^{\text{II}}$ $(K=17160)$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ $(K=85800)$, where $K\!=\!k\cdot L\cdot Z$. The maximum frame lengths of $\mathcal{C}_{\text{SC}}^{\text{I}}$$\mathcal{C}_{\text{SC}}^{\text{II}}$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ are $N=25311$$N=50622$, and $N=253110$ for the lowest rate of $R=0.34$, respectively, which are approximately 1×1$\times$1 ×2×2$\times$2 ×, and 10×10$\times$10 × longer compared to 5G-NR LDPC codes with the same rate. [^5] Note that in this paper, we only consider the maximum blocklength 5G-NR LDPC codes with $Z=384$ to fully present the 5G-NR performance. In SLME decoding, we let $W=S=L+\omega=11$, thus decoding is configured to run BP on the entire $\mathbf{H}_{\text{SC}}$ of ESRL codes. Given that hardware implementations as in [22] and [48] typically limit the iteration count to $5$ or below to minimize decoding delay and enhance throughput, in Fig. 11, we set $I_{\text{max}}=5$ to highlight the ESRL performance in these high-throughput scenarios. At $R=0.87$, since the ESRL code only uses the HRC part without involving the IME, a slight error floor appears at the FER of $10^{-5}$. At other rates, ESRL codes exhibit a FER performance without error floors and achieve stable performance gains when $Z$ increases. For instance, $\mathcal{C}_{\text{SC}}^{\text{III}}$ with $Z=390$ outperforms $\mathcal{C}_{\text{SC}}^{\text{I}}$ with $Z=39$ by $0.35$ dB when $R=0.72$.

### v2

@@ IV-A2 Spreading Optimization over the Unified Graph
Unlike [33, 30], there is no need to repeatedly run the above message-passing algorithm on $\mathcal{G}_{\text{U}}$ to update the cycle distribution when changes occur in $\mathcal{G}_{\text{U}}$. Instead, we can simply adjust the polynomial $p_{v}$ in (14) to measure the impact of reallocating a specific edge on the cycle distribution. It is noteworthy that the extra complexity of such an optimization on $\mathcal{G}_{\text{U}}$ is negligible for a small $\omega$, such as the frequent choice of $\omega=1$. Specifically, if the edge $e_{v}^{c}$ is re-assigned from $\mathbf{T}_{c,v}$ to ${\widetilde{\mathbf{T}}_{c,v}}$ (where ${\widetilde{\mathbf{T}}_{c,v}}\in[0,\omega]$ and ${\widetilde{\mathbf{T}}_{c,v}}\neq\mathbf{T}_{c,v}$), let $\theta={\widetilde{\mathbf{T}}_{c,v}}-\mathbf{T}_{c,v}$, we modify the original $p_{v}$ by two steps: 1 Add $\theta$ to all accumulated path values $\phi$ of the initial $\gamma_{c}$ variable in the polynomial $p_{v}$ 2 Subtract $\theta$ from the $\phi$ of all $\gamma$ variables in the monomial passed to the VN $v$ along the edge $e_{c}^{v}$ (as both $e_{v}^{c}$ and $e_{c}^{v}$ belong to the same undirected edge, they are equivalent in the spreading process).
@@ 5       Update $\mathcal{G}_{\text{U}}$ by $\widetilde{\mathbf{T}}_{c,v}^{\ast}$;
IV-A3 Computational Complexity
Following the extrinsic message-passing principle, the computational complexity of our algorithm is primarily attributed to the update of accumulated path values $\phi$ according to (12) and (13). For $l$-cycles, the total complexity of cycle counting over $\mathcal{G}_{\text{U}}$ is $O(l\cdot|\mathbb{E}|\cdot|\mathcal{M}|)$, where $|\mathcal{M}|$ represents the average number of $\gamma$ variables in a monomial. This value is dynamic and upper bound by $\lceil 2Q_{v}^{l}/d_{v}\rceil$. To eliminate $l$-cycles on $\mathcal{G}_{\text{U}}$, the overall complexity of Algorithm 1 is $O(I_{\text{MP}}\cdot\omega\cdot|\mathbb{E}|\cdot|\mathcal{M}|)$ when $\omega$ is large. However, if $\omega\leq l$, this complexity will reduce to $O(I_{\text{MP}}\cdot\l\cdot|\mathbb{E}|\cdot|\mathcal{M}|)$. Namely, for a small $\omega$, the additional complexity (line $4$ of Algorithm 1) introduced by this optimization on $\mathcal{G}_{\text{U}}$ becomes negligible compared to the cycle counting part.
@@ 6 Step 3: Optimize over unified graph $\mathcal{G}_{\text{U}}$
7    $\mathbf{T}\leftarrow\mathbf{0}$; //initialization
@@ 14 Repeat Step 2 to Step 4 $I_{\text{HRC}}$ times;
The HRC design process is outlined in Algorithm 2. In Step 1, under the ESRL code constraints from Section III-B, DE is used to determine the optimal distribution of $d_{v}$, denoted as the polynomial $p_{d_{v}}^{\ast}$. Step 2 employs the progressive edge growth (PEG) algorithm [55] and the approximate cycle extrinsic message degree (ACE) constraint [56] to maximize local girth and to ensure the graph connectivity of short cycles on $\mathbf{B}_{\text{HRC}}$, respectively, which are two classical methods for designing finite-length LDPC codes. Notably, threshold analysis on $\mathbf{B}_{\text{HRC}}$ is not part of our flow, aligning with the objective of ESRL codes in Section III-A to focus on constructing a better coupled protomatrix $\mathbf{B}_{\text{SC}}$. The next step is to use the unified graph $\mathcal{G}_{\text{U}}$ in Section IV-A and the modified coupled RCA in Section IV-B to optimize $\mathbf{B}_{\text{SC}}$. In Step 3, the matrix $\mathbf{T}$ is initially set to $\mathbf{0}$, indicating that all edges are initially assigned to $\mathbf{B}_{0}$. The unified graph $\mathcal{G}_{\text{U}}$ is then employed to optimize the spreading strategy, reallocating each edge to minimize cycle count. In Step 4, $\mathbf{B}_{\text{SC}}$ is constructed using the code profile $\mathbf{B}$$\mathbf{T}$, and $\mathbf{Q}$, followed by modified coupled RCA to compute the threshold. Since $\mathbf{Q}$ has a fixed structure without additional design, it directly participates in the threshold analysis of $\mathbf{B}_{\text{SC}}$. Finally, after $I_{\text{HRC}}$ iterations, the optimal $\mathbf{B}_{\text{HRC}}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold are selected.
The IME design process is summarized in Algorithm 3. Compared to the overall design of the HRC in Algorithm 2, the IME is constructed row-wise until the full matrix of size $m\times n$ is completed. In the ESRL code, we find that the performance of the growing IME matrix is highly sensitive to the previously designed sections, especially after edge spreading, i.e., overemphasizing the performance of a specific intermediate rate will lead to significant performance degradation in subsequent matrix extensions. To address this issue, instead of using a greedy approach to explore a local optimum for each newly added row as in [18], we start with a function called ${\small\textsf{GlobalWeightRef}}(\cdot)$ in Step 1, which essentially runs Algorithm 2 to design a reference $\mathbf{B}$ of size $m\times n$ directly and thus offers a preliminary IME example. Note that although this reference is not optimal, it can provide a rough weight guideline $d_{\text{IME}}$ for the IME design to avoid being trapped in local optima for intermediate rows. Similarly, Step 2 involves using the PEG and ACE algorithms for the newly added row in $\mathbf{T}$, which is initially set to zero. In Step 2.2, the unified graph $\mathcal{G}_{\text{U}}$ is generated to optimize the spreading strategy. It is noteworthy that when designing the $c$-th row of the IME, the edges and corresponding allocations of previous $c-1$ rows must remain unchanged to preserve the optimization results obtained before. Besides, certain special structures in ESRL codes, such as assigning the diagonal of the SSC to $\mathbf{B}_{0}$, need to be retained. Finally, we run the process $I_{\text{IME}}$-times for each row to search for optimal $\mathbf{B}^{\ast}$ and $\mathbf{T}^{\ast}$ with the lowest threshold, until completing the entire ESRL protomatrix design $\mathbf{B}_{\text{SC}}$.
@@ 2    $\mathbf{B}^{\ast}\leftarrow\mathbf{B}_{\text{HRC}}^{\ast}$;
4    $d_{\text{IME}}\leftarrow{\small\textsf{GlobalWeightRef}}(\mathbf{B}_{\text{HRC}},m,n)$;
@@ 5 Step 2: Generate IME row by row
8          $\mathbf{B}\leftarrow\mathsf{PEG}(\mathbf{B}^{\ast},d_{\text{IME}})$;
@@ V Puncturing Scheme and Decoding Algorithm
In this section, we introduce the puncturing scheme, describe the integration in the IR-HARQ process, and explain high-throughput SLME decoding tailored to ESRL codes.
Figure 8: Puncturing scheme of the ESRL code, where the proposed semi-layered multi-engine decoding is running [57].
@@ V-A Puncturing Scheme
V-B IR-HARQ with ESRL Codes
We now demonstrate that ESRL codes can support the implementation of IR-HARQ. When decoding an ESRL codeword $\bm{c}_{[0,L-1]}$ fails, where $\bm{c}_{[i]}\triangleq[\bm{u}_{[i]},\bm{r}_{[i]}]$, sending $L$ batches of additional parity redundancy $\bm{r}_{{[i]}_{\Delta}}$ can still form a valid codeword of the ESRL code at a lower code rate. The incremental codeword of each batch, $\bm{c}_{[i]}^{\prime}=[\bm{u}_{[i]},\bm{r}_{[i]}^{\prime}]=[\bm{u}_{[i]},\bm{r}_{[i]},\bm{r}_{{[i]}_{\Delta}}]$, combines to create $\bm{c}_{[0,L-1]}^{\prime}=[\bm{c}_{[0]}^{\prime},\bm{c}_{[1]}^{\prime},\dots,\bm{c}_{[L-1]}^{\prime}]$, as shown in Fig. 9.
Figure 9: IR-HARQ with ESRL codes.
Proof.
For brevity, we restrict $\omega$ of ESRL codes to $1$, while cases with $\omega>1$ can be derived in a similar way. Based on SC-LDPC encoding [46]$\bm{r}_{[i]}=\widetilde{\mathbf{H}}_{0}\cdot\bm{u}_{[i]}+\mathbf{H}_{1}\cdot\bm{c}_{[i-1]},i\geq 1$, where $\mathbf{H}_{0}$ is transformed as $\mathbf{H}_{0}\triangleq[\widetilde{\mathbf{H}}_{0},\mathbf{I}]$. We assume the $(i-1)$-th batch has already satisfied rate compatibility, i.e., $\bm{c}_{[i-1]}^{\prime}=[\bm{u}_{[i-1]},\bm{r}_{[i-1]}^{\prime}]=[\bm{u}_{[i-1]},\bm{r}_{[i-1]},\bm{r}_{{[i-1]}_{\Delta}}]$. Hence, in combination with Fig. 9, we can obtain
$\mathbf{H}_{1}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}=[\mathbf{H}_{1}\cdot\bm{c}_{[i-1]};\mathbf{H}_{{1}_{\Delta}}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}].$ (22)
Since $\widetilde{\mathbf{H}}_{0}^{\prime}=[\widetilde{\mathbf{H}}_{0};\widetilde{\mathbf{H}}_{{0}_{\Delta}}^{\prime}]$ in Fig. 9, we easily derive $\widetilde{\mathbf{H}}_{0}^{\prime}\cdot\bm{u}_{[i]}=[\widetilde{\mathbf{H}}_{0}\cdot\bm{u}_{[i]};\widetilde{\mathbf{H}}_{{0}_{\Delta}}^{\prime}\cdot\bm{u}_{[i]}]$. Hence, $\bm{r}_{[i]}^{\prime}$ of the $i$-th batch after incremental redundancy transmission is computed as (23)
$\bm{r}_{[i]}^{\prime}$ $=\widetilde{\mathbf{H}}_{0}^{\prime}\cdot\bm{u}_{[i]}+\mathbf{H}_{1}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}$ (23)
$=[\widetilde{\mathbf{H}}_{0}\cdot\bm{u}_{[i]}+\mathbf{H}_{1}\cdot c_{[i-1]};\widetilde{\mathbf{H}}_{0_{\Delta}}^{\prime}\cdot\bm{u}_{[i]}+\mathbf{H}_{1_{\Delta}}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}]$
$=[\bm{r}_{[i]};\widetilde{\mathbf{H}}_{0_{\Delta}}^{\prime}\cdot\bm{u}_{[i]}+\mathbf{H}_{1_{\Delta}}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}].$
Herein, we can derive $\bm{r}_{[i]}^{\prime}=[\bm{r}_{[i]},\bm{r}_{{[i]}_{\Delta}}]$ if $\bm{r}_{{[i]}_{\Delta}}=\widetilde{\mathbf{H}}_{0_{\Delta}}^{\prime}\cdot\bm{u}_{[i]}+\mathbf{H}_{1_{\Delta}}^{\prime}\cdot\bm{c}_{[i-1]}^{\prime}$. Subsequently, this property can be recursively proven for all spatial positions $i$$0\leq i<L$. As shown in (24), when decoding of $\bm{c}_{[0,L-1]}$ fails, the incremental codeword $\bm{c}_{[0,L-1]}^{\prime}$ remains valid if only sending additional parity redundancy bits.
$\bm{c}_{[0,L-1]}^{\prime}\!=\![\bm{u}_{[0]},\underbrace{\bm{r}_{[0]},\bm{r}_{[0]_{\Delta}}}_{\bm{r}_{[0]}^{\prime}},\ldots,\!\bm{u}_{[L-1]},\underbrace{\bm{r}_{[L-1]},\bm{r}_{{[L-1]}_{\Delta}}}_{\bm{r}_{[L-1]}^{\prime}}],$ (24)
[…]
$\quad\quad\quad\quad\bm{c}_{[0,L-1]}\!=\![\bm{u}_{[0]},\bm{r}_{[0]},\ldots,\bm{u}_{[L-1]},\bm{r}_{[L-1]}].$
A comparison between ESRL codes and 5G-NR LDPC codes with respect to the performance under IR-HARQ is provided in Section VI-D.
V-C SLME Decoding of ESRL codes
@@
In this section, we propose a high-throughput semi-layered decoding algorithm using a multi-engine architecture in [13], tailored to ESRL codes. This algorithm can be flexibly configured as either high-performance full BP decoding of the entire $\mathbf{H}_{\text{SC}}$ or as low-complexity windowed BP sliding along the diagonal of $\mathbf{H}_{\text{SC}}$. Let $W$ denote the window size, $S$ the sliding step, and $I_{\text{max}}$ the maximum number of decoding iterations within the window at each step. For instance, by setting $W\!=\!S\!=\!L+\omega$, SLME decoding corresponds to full BP decoding with $I_{\text{max}}$ for ESRL codes.
@@
While coupling generally limits parallel decoding, ESRL codes without parallel edges can perform semi-layered decoding through a more compact multi-engine architecture compared to [13], achieving higher parallelism. Specifically, we apply individual layered decoding to each $\mathbf{U}$ (defined in (3) of Section III-A and also shown in Fig. 3), without causing memory read-write conflicts. Two adjacent engines overlap and share the posterior LLR memory of $\omega$ batches. Thanks to the overlapping between engines, even with parallel processing across engines, each engine can still utilize the updated LLRs of the previous $\omega$ batches provided by the preceding engines during layered decoding. The ‘semi-layered’ nature of SLME decoding therefore preserves the important data dependencies of ESRL codes, while the ‘multi-engine’ approach significantly enhances decoding parallelism. In Fig. 8, with a window size of $W=4$, conventional layered decoding proceeds row-wise from top to bottom (indicated by a grey arrow). In contrast, SLME decoding can deploy four identical sub-decoders working in parallel (each running layered decoding shown by a light red arrow in Fig. 8), thus resulting in a 4×4$\times$4 × throughput gain in the example.
@@ VI Design Example and Performance Comparisons
Figure 11: ACE spectrum of the protograph example of ESRL codes.
@@ VI-A Design Example of ESRL Codes
Fig. 10 illustrates an example ESRL protomatrix family designed for $k=22$ (in each batch). We intentionally set $k$ equal to that of 5G-NR to facilitate fair comparisons in Section VI-B. Following SC-LDPC decoder implementations [50, 58] to balance performance and complexity, we set $\omega=1$ to simplify the code design. Moreover, the coupling length $L$ is set as $10$ to balance the rate loss and the underlying blocklength for a given frame length. Let $m^{\prime}=4$$n^{\prime}=26$$m=40$, and $n=62$. Hence, the HRC protomatrix has dimensions of $4\times 26$. The IME consists of the IRC protomatrix (sized $36\times 26$) and the SSC protomatrix (sized $36\times 36$). Considering that the first VN in the HRC is always punctured ($\rho=1$), the example ESRL code can support $37$ distinct rates, ranging from the highest rate of $0.87$ to the lowest rate of $0.34$. The uncoupled protomatrix $\mathbf{B}$ has a size of $40\times 62$ and a weight of $314$. Note that with this configuration we mark six classical rates in Fig. 10: $0.87~{}(4\times 26)$$0.8~{}(6\times 28)$$0.72~{}(9\times 31)$$0.63~{}(13\times 35)$$0.52~{}(20\times 42)$, and $0.34~{}(40\times 62)$, which are also supported by 5G-NR and are used for the subsequent comparisons. For instance, the highest rate is computed as $R=(22\times 10)/(26\times 10+4-1\times(10+1))\approx 0.87$ using (6). It is noteworthy that since the HRC is relatively dense, for $R=0.87$, there exist some elementary trapping sets with the number of VNs as $2$ or $3$, which explains the onset of the error floors in Fig. 12 and Fig. 13. The edge-spreading matrix $\mathbf{T}$ is also reported in Fig. 10, with $\mathbf{B}_{0}$ and $\mathbf{B}_{1}$ distinguished by colors. As mentioned in Section III-C, the tail matrix $\mathbf{Q}$ has a fixed tri-diagonal structure, marked as black in Fig. 10. In terms of connectivity of short cycles, in the lowest rate protograph, every cycle of length up to $8$ is connected to at least $12$ extrinsic CNs. For $4$-cycles, the value of ACE $\eta$ ranges from $6$ to $12$. Fig. 11 also shows the ACE spectrum of $6$-cycles and $8$-cycles. It is noteworthy that given these ACE values are obtained from the protograph before lifting, the actual graph connectivity will be even better in $\mathbf{H}_{\text{SC}}$ of the ESRL code after lifting.
@@
Fig. 12 shows the FER performance of the example ESRL codes under different lifting sizes. Let $\mathcal{C}$ define a series of rate-compatible LDPC codes with $K$ information bits, and $\mathcal{C}(N,K)$ be one of its code patterns with a rate $R=K/N$. Based on $Z\in\{39,78,390\}$, three series of ESRL codes are defined: $\mathcal{C}_{\text{SC}}^{\text{I}}$ $(K=8580)$$\mathcal{C}_{\text{SC}}^{\text{II}}$ $(K=17160)$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ $(K=85800)$, where $K\!=\!k\cdot L\cdot Z$. The maximum frame lengths of $\mathcal{C}_{\text{SC}}^{\text{I}}$$\mathcal{C}_{\text{SC}}^{\text{II}}$, and $\mathcal{C}_{\text{SC}}^{\text{III}}$ are $N=25311$$N=50622$, and $N=253110$ for the lowest rate of $R=0.34$, respectively, which are approximately 1×1$\times$1 ×2×2$\times$2 ×, and 10×10$\times$10 × longer compared to 5G-NR LDPC codes with the same rate. [^5] Note that in this paper, we only consider the maximum blocklength 5G-NR LDPC codes with $Z=384$ to fully present the 5G-NR performance. In SLME decoding, we let $W=S=L+\omega=11$, thus decoding is configured to run BP on the entire $\mathbf{H}_{\text{SC}}$ of ESRL codes. Given that hardware implementations as in [27, 59, 60] typically limit the iteration count to $5$ or below to minimize decoding delay and enhance throughput, we set $I_{\text{max}}=5$ to highlight the ESRL performance in these high-throughput scenarios, as shown in Fig. 12. At $R=0.87$, since the ESRL code only uses the HRC part without involving the IME, a slight error floor appears at the FER of $10^{-5}$. At other rates, ESRL codes exhibit a FER performance without error floors and achieve stable performance gains when $Z$ increases. For instance, $\mathcal{C}_{\text{SC}}^{\text{III}}$ with $Z=390$ outperforms $\mathcal{C}_{\text{SC}}^{\text{I}}$ with $Z=39$ by $0.35$ dB when $R=0.72$ at the FER of $10^{-5}$.
Figure 14: FER comparisons of ESRL codes and 5G-NR LDPC codes [4] at $R=0.8$ with different $I_{\text{max}}$.